12x-7 5x+18
______ = _______
4 3

Answer by checkley77(12844)

12x-7 5x+18
______ = _______ cross multiply.
4 3
3(12x-7)=4(5x+18)
36x-21=20x+72
36x-20x=72+21
16x=93
x=93/16 ans.
3(12*93/16-7)=4(5*93/16+18)
3(1116/16-7*160/16=4(465+18*16)/16
3(1116-112)/16=4(465+288)/16
3*1004/16=4*753/16
3012/16=3012/16